If i use 0.14 ml per day, how many days will it take me to use 100 ml?
Answer by TimothyLamb(4379)

r = v/t
r = 0.14/1
r = 0.14
---
t = v/r
t = 100/0.14
t = 714.285714286 days
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---